A large group of students took a test in Physics and the final grades have a mean of 70 and a standard deviation of 10. If we can approximate the distribution of these grades by a normal distribution, what percent of the students

  i) scored higher than 80?                                                                                                                                          (3mks)

  ii) should pass the test?                                                                                                                                           (3mks)


1
Expert's answer
2021-02-24T06:02:38-0500

Let X be random variable of grades a random student scores.

X is normally distributed with a mean "\mu = 70" and a standard deviation "\sigma = 10". Therefore, "X = \sigma Y +\mu" where random variable Y has a standard Gaussian distribution with the cumulative distribution function "P(Y<y) =\Phi(y) = \frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{y}e^{-t^2/2}dt"

"P(x\leq X) = P(x\leq \sigma Y +\mu) = P(\frac{x-\mu}{\sigma} \leq Y) = 1 - \Phi(\frac{x-\mu}{\sigma})"


Let "x=80" then "(x-\mu)/\sigma = (80 - 70)/10 = 1"

"P(80\leq X) = 1 - \Phi(1)= 1-0.84134 = 0.15866=15.866\%"

Therefore, approximately 15.866% of students score 80+.


A student passes the test if he scores 60+.

Let "x=60" then "(x-\mu)/\sigma = (60 - 70)/10 = -1"

"P(60\leq X) = 1 - \Phi(-1)= 1-0.15866=0.84134"

Therefore, approximately 84.134% of students pass the test on Physics.


Answer.

Approximately 15.866% of students score 80+

Approximately 84.134% of students pass the test